# GTPBD: A Fine-Grained Global Terraced Parcel and Boundary Dataset

### v1

@@ 4.4 Edge Detection Results
Conventional edge detection datasets typically include multi-scale, manually annotated boundaries. To simulate this for the GTPBD dataset, we apply morphological operations (erosion and dilation) to generate edge labels with widths of 1 to 5 pixels, mimicking human-labeled multi-resolution boundaries. Table 3 presents the results of four edge detection models. Among them, REAUNet-Sober achieves the best performance across all three metrics—ODS (0.6506), OIS (0.6373), and AP (0.7009)—significantly outperforming both general-purpose (e.g., PiDiNet), demonstrating its effectiveness in delineating parcel edges under noisy conditions. Visual examples in Fig. 7 (f) further illustrate model robustness, where false positives are marked in red and missed detections in blue.
@@
U-Net 83.27 75.69 65.22 79.87 78.95 66.11 37.33 38.80
PSPNet 83.70 80.56 69.64 82.23 82.10 64.75 38.75 40.15
SegFormer 82.45 82.19 69.95 82.24 82.32 63.39 39.85 40.87
Mask2Former 82.59 81.83 69.80 82.19 82.21 62.98 39.40 40.55
REAUNEet 78.44 76.79 63.41 78.98 77.61 23.77 53.05 25.18
SEANet 69.85 88.58 64.08 76.54 78.11 11.50 71.96 14.82
HGBNet 76.14 87.82 68.87 81.17 81.56 8.58 67.39 12.22
@@
S $\rightarrow$ N Source Only 75.93 82.12 70.05 77.47 78.92 73.11 83.21 68.36 76.71 77.87 8.94 49.80 11.68
FDAFDA 82.60 78.20 67.13 79.27 80.34 79.55 77.95 64.94 78.26 78.74 25.04 46.09 24.44
PiPa PiPa 88.40 81.43 73.57 84.16 84.77 88.40 81.43 73.57 84.16 84.77 24.59 16.63 11.90
HRDA HRDA 88.00 82.44 74.11 84.48 85.13 84.00 83.44 72.00 83.24 83.72 18.89 41.22 19.87
DAFormer DAFormer 87.23 82.71 73.78 84.08 84.91 84.83 83.63 72.75 83.72 84.23 17.52 39.94 19.22
S $\rightarrow$ G Source Only 74.66 68.19 60.69 72.16 75.54 71.54 70.14 60.53 72.44 75.41 24.59 49.02 23.90
FDAFDA 76.27 87.16 68.57 74.78 81.35 81.18 75.68 63.38 74.78 78.33 22.11 48.85 22.81
PiPa PiPa 83.90 75.22 65.73 75.25 79.32 83.90 75.22 65.73 75.25 79.32 29.28 21.09 12.94
HRDA HRDA 61.35 95.53 59.64 66.31 74.71 78.22 77.30 63.61 73.36 77.76 20.93 50.11 21.94
DAFormer DAFormer 82.89 80.27 68.87 77.10 81.56 79.64 82.05 67.82 76.55 80.82 17.53 49.80 21.01
N $\rightarrow$ S Source Only 64.35 80.32 55.58 65.93 61.45 54.97 78.98 47.95 63.27 64.82 13.83 73.75 19.95
FDAFDA 67.40 71.73 53.25 67.20 69.50 59.08 68.04 46.25 66.11 63.25 29.48 60.68 33.65
PiPa PiPa 77.68 61.11 51.98 70.59 68.40 77.68 61.11 51.98 70.59 68.40 42.26 30.86 26.13
HRDA HRDA 76.90 60.10 50.91 69.81 67.47 66.52 65.74 49.39 71.14 66.13 32.03 58.84 36.01
DAFormer DAFormer 79.64 60.69 52.54 71.44 68.89 68.68 68.03 51.92 73.01 68.35 25.72 62.58 33.00
N $\rightarrow$ G Source Only 72.94 72.51 63.10 73.27 77.38 70.05 74.88 63.10 73.62 77.38 21.50 51.35 22.35
FDAFDA 84.17 74.95 65.69 75.30 79.29 81.18 75.68 63.38 74.78 78.33 22.11 48.85 22.81
PiPa PiPa 83.52 76.03 66.11 75.40 79.60 83.52 76.03 66.11 75.40 79.60 27.74 22.01 13.36
HRDA HRDA 87.45 66.81 60.96 73.01 75.75 84.59 68.04 60.54 73.28 75.42 27.13 42.88 24.07
DAFormer DAFormer 83.56 77.49 67.24 76.17 80.41 79.98 78.75 65.79 75.32 79.36 18.40 49.54 21.72
G $\rightarrow$ N Source Only 76.82 80.44 69.52 77.30 78.59 74.21 81.99 68.46 76.96 77.93 9.71 48.20 12.46
FDAFDA 71.77 89.22 66.04 75.16 79.55 72.10 82.14 62.33 74.35 76.79 21.83 54.51 24.76
PiPa PiPa 74.40 92.08 69.93 78.56 82.30 74.40 92.08 69.93 78.56 82.30 10.00 34.23 9.47
HRDA HRDA 75.42 91.37 70.41 79.20 82.63 72.67 90.95 67.77 77.66 80.79 9.33 52.11 13.88
DAFormer DAFormer 85.40 79.72 70.16 81.63 82.46 81.99 79.93 67.99 80.56 80.95 19.52 44.51 20.39
G $\rightarrow$ S Source Only 54.12 71.52 45.98 55.97 61.78 54.65 80.53 48.27 63.02 65.11 13.66 74.62 20.22
FDAFDA 59.39 88.36 55.09 62.47 71.04 51.21 81.85 45.99 58.81 63.01 21.59 66.56 28.27
PiPa PiPa 65.25 88.79 60.29 69.53 75.22 65.25 88.79 60.29 69.53 75.22 16.14 32.04 12.58
HRDA HRDA 61.35 95.53 59.64 66.31 74.71 66.30 95.09 60.33 70.05 75.33 16.20 66.58 28.68
DAFormer DAFormer 70.04 91.15 63.72 70.60 77.11 67.77 91.66 61.91 70.65 75.18 14.32 64.13 26.83
Avg Source Only 69.14 75.52 60.49 70.01 72.61 66.43 76.82 59.11 69.52 73.76 15.87 56.93 18.10
FDAFDA 73.60 81.60 61.96 72.53 77.18 70.55 76.89 58.38 72.38 75.38 23.36 54.09 25.46
PiPa PiPa 78.86 79.77 64.94 75.91 78.27 80.06 77.78 65.61 77.87 78.47 24.00 26.81 14.73
HRDA HRDA 78.18 81.05 62.28 74.52 78.40 75.72 77.43 60.86 76.79 76.53 24.12 51.79 24.41
DAFormer DAFormer 81.13 78.67 66.39 76.84 79.56 77.15 77.84 64.82 76.63 78.16 19.58 53.90 23.36
@@ 4.5 Boundary-integration terraced parcel extraction results
Building on the boundary constraint theory by Yuan et al. Yuhui2020 , which emphasizes the role of edge orientation and intensity in enhancing object-level delineation, we explore boundary-integration strategies for agricultural parcel segmentation. We compare four general-purpose segmentation models (U-Net, PSPNet, SegFormer, Mask2Former) with three agriculture-specific ones (HBGNet FHAPD/HBGNet , REAUNet ReauNet-Sober , SEANet SEANet ). To incorporate edge cues, we introduce a boundary detection pipeline. REAUNet-Sober ReauNet-Sober achieves the best performance among trained edge detectors, and its binarized output is integrated into the segmentation process. As shown in Table 4, REAUNet ReauNet-Sober delivers the best object-level performance with the lowest GUC with 53.05%, despite lower pixel-level scores. In contrast, SegFormer segformer attains the highest pixel-level accuracy (OA: 82.24%) but suffers from higher object-level errors, indicating a trade-off between pixel precision and structural integrity.
@@ 4.6 Unsupervised domain adaptation results
We evaluate domain robustness of segmentation models through UDA experiments across three distinct domains in GTPBD (see Sec. 3.3). The Source only and other four domain adaptation methods are benchmarked: FDA, DAFormer, HRDA, and PiPa. As shown in Table 5, under standard (non-boundary-integration) settings, DAFormer DAFormer achieves the highest average performance across Prec., IoU, OA, and F1, demonstrating strong generalization under domain shift. After adding boundary results (integrated by the performance of REAUNet-Sober ReauNet-Sober ), PiPa PiPa consistently outperforms others, achieving top scores across both pixel-level (Prec., IoU, OA, F1) and object-level metrics, including the lowest GUC and GTC. This highlights its advantage in preserving parcel structure with fine-grained segmentation. Visualization comparisons are provided in Appendix E.4.
@@ 5 Conclusion
Appendix A Broader impact
Appendix B More image cases
B.1 More image cases in GTPBD
Figure 8: More images cases of different regions in GTPBD.
GTPBD covers seven major geographic zones in China and transcontinental climatic regions around the world. There are different terraces in different regions, and Fig. 8 supplements the images of terraces of the 12 regions mentioned in Fig. 2 .
B.2 Some cases in previous datasets
Fig. 9 displays typical annotation–imagery pairs from six widely used parcel datasets: (a) FHAPD [48], (b) FTW [17], (c) AI4Boundaries [9], (d) PASTIS [12], (e) CP-Set [46], and (f) GFSAD30 [31]. Although each benchmark has advanced parcel delineation research, several common limitations are evident. FHAPD and CP-Set rely on binary masks with coarse edges, leaving thin ridges either over-smoothed or missing. FTW and PASTIS are built on Sentinel-2 imagery (10 m GSD), so images appear blurry and smallholder fields are merged. AI4Boundaries depicts large, orthogonal parcels from mechanised farms, providing little shape diversity. GFSAD30, at 30 m resolution, captures only blocky field outlines and omits fine-scale geometry altogether. These issues—low spatial resolution, overly regular parcel shapes, and imprecise or incomplete boundaries limit the evaluation of models on fragmented, irregular terrains.
B.3 More statistics of GTPBD
Appendix C Model details
C.1 Semantic segmentation
C.1.1 U-Net
U-Net, proposed by Ronneberger et al. [28], adopts a symmetric encoder–decoder design that captures contextual information via a contracting path and enables precise localization through an expansive path. In our experiments using MMSegmentation’s default U-Net implementation, the key hyperparameters are configured as Table 6.
Table 6: The hyperparameters of U-Net
Hyperparameter Description Typical Value Input image size Dimensions of input images 512×512 Backbone base channels Number of filters in the first convolution layer 64 Number of stages Levels of encoder–decoder (depth of U-Net) 5 Convolutions per stage Number of conv layers in each block 2 Optimizer Optimization algorithm and initial learning rate Adam, lr=$1e^{-4}$ Sliding-window crop size Sliding-window inference patch size 256 × 256 Sliding-window stride Stride between adjacent sliding-window patches 170 Training iterations Total number of training iterations 20,000
C.1.2 DeepLabV3
DeepLabV3 [6] enhances semantic segmentation by employing atrous (dilated) convolutions to enlarge the receptive field without losing resolution, and by integrating an Atrous Spatial Pyramid Pooling (ASPP) module to capture multi-scale context. In our experiment, All hyper-parameters remain the same as in the mmsegmentation framework and are summarized in Table 7.
Table 7: The hyperparameters of DeeplabV3
C.1.3 PSPNet
PSPNet [49] introduces a Pyramid Pooling Module that aggregates context information at multiple spatial scales by pooling features into different bin sizes, and then upsampling and concatenating them with the original feature map. This design enables the network to capture both global context and fine details simultaneously. In our experiments, we rely on the out-of-the-box configuration provided by MMSegmentation without further modification, and summarize the key hyperparameters in Table 8.
Table 8: The hyperparameters of PSPNet
C.1.4 NonLocal
Non-local Neural Networks [38], a generic non-local operation that computes pairwise interactions between any two positions on the feature map, enabling the model to capture long-range dependencies and context beyond local receptive fields. We adopt the standard MMSegmentation implementation without modifying its default settings and report the core hyperparameters in Table 9.
Table 9: The hyperparameters of NonLocal
C.1.5 OCRNet
OCRNet [43] enhances pixel-wise segmentation by introducing an object-contextual representation module that aggregates contextual information from object regions to refine per-pixel predictions. Specifically, it first generates a coarse object region map via soft object-region pooling, then computes a pixel-wise contextual representation by weighting features according to their belongingness to these regions. This two-stage design allows the network to capture both local details and global object semantics effectively. Table 10 summarize the main hyperparameters in our experiment.
Table 10: The hyperparameters of OCRNet
C.1.6 K-Net
K-Net [47] formulates semantic segmentation as a dynamic kernel-based grouping problem, enabling adaptive context aggregation and object-level reasoning. Each kernel generates an attention map that segments the feature map into distinct semantic regions, and these kernels are updated via a lightweight transformer-style interaction. In our experiments, we employ MMSegmentation’s K-Net configuration without further modifications and summarize the core hyperparameters in Table 11.
Table 11: The hyperparameters of K-Net
C.1.7 Segformer
SegFormer [39] leverages a hierarchy of lightweight MLP-based Mix-FFN blocks and multi-level feature fusion to achieve efficient and effective semantic segmentation without positional encodings. Its simple encoder–decoder design employs a series of Transformer-like layers that progressively downsample the input, and fuse multi-scale representations via a lightweight MLP decoder. Key parameters are shown in Table 12 in our experiment.
Table 12: The hyperparameters of Segformer
C.1.8 Mask2Former
Mask2Former [8] presents a universal architecture for segmentation tasks by modeling pixel-wise, instance-wise, and semantic-level grouping using a mask-based Transformer decoder. It decouples mask prediction from task-specific heads, employing dynamic queries and multi-scale features to produce segmentation masks across different granularities. We employ Mask2Former configuration directly without any additional tuning and the key hyperparameters can be seen in Table 13.
Table 13: The hyperparameters of mask2former
C.2 Edge Detection Methods
C.2.1 UEAD
UEAD [53] addresses the annotation ambiguity and subjectivity inherent in edge detection by modeling label uncertainty. Instead of relying on deterministic pixel-wise labels, it learns a Gaussian distribution over annotations to capture labeling variance. The estimated variance serves as a measure of pixel-level uncertainty, and an adaptive weighting loss emphasizes learning from uncertain (i.e., hard) samples. UEAD can be integrated with various encoder–decoder backbones and consistently improves performance across multiple edge detection benchmarks.
C.2.2 MuGE
MuGE[52] introduces a novel edge detection framework that captures edge ambiguity by generating edge maps at multiple controllable granularities. It first predicts edge granularity from annotations, then injects this granularity into multi-scale feature maps to produce edge maps ranging from coarse contours to fine textures. By decomposing feature maps into frequency components, MuGE enables fine control over edge detail, resulting in both interpretability and high accuracy.
C.2.3 PiDiNet
PiDiNet[30] offers a lightweight and efficient solution for edge detection by combining traditional gradient operators with modern convolutional design. It introduces pixel difference convolutions that emulate classical detectors like Canny and Sobel, enabling real-time inference with high accuracy. With less than 1M parameters, PiDiNet achieves human-level performance on BSDS500 and maintains strong efficiency–accuracy trade-offs across diverse benchmarks.
C.3 Agricultural Parcel Extraction Methods
C.3.1 REAUNet
REAUNet[21] is a tailored edge-aware network designed for accurate agricultural parcel delineation from both medium- and high-resolution remote sensing imagery (Sentinel-2 and GF-2). It addresses the common issue of unclosed and fragmented parcel edges by integrating four key components: an edge detection block, a dual attention block, a deep supervision mechanism, and a refine module. These modules collectively enhance the model’s sensitivity to boundary information and improve multi-scale consistency. Experiments show that all four components are critical, with full REAUNet outperforming partial variants. Compared with SEANet and MPSPNet, REAUNet achieves improvements in both thematic (F1, IoU) and geometric (GTC) accuracy, and demonstrates promising transferability to unseen regions with limited fine-tuning data.
C.3.2 SEANet
SEANet[18] is a multi-task neural network that simultaneously predicts semantic masks, edges, and distance maps for precise and closed parcel delineation. By explicitly modeling boundary extraction as an edge detection task, SEANet enhances geometric accuracy and is particularly effective for small, irregular parcels. It incorporates a multi-level edge feature extraction mechanism and a task uncertainty-aware loss to improve generalization. Extensive experiments on both high-resolution (GF-2) and medium-resolution (Sentinel-2) images across China and Europe show that SEANet produces more accurate parcel layouts and demonstrates robust cross-region transferability.
C.3.3 HBGNet
HBGNet[48] is a hierarchical dual-branch framework designed to fully exploit boundary semantics for robust agricultural parcel extraction. It consists of a core AP extraction branch and an auxiliary boundary branch enhanced by Laplacian-based convolution. To improve adaptability across varying parcel sizes and morphologies, HBGNet integrates global–local context aggregation and boundary-guided fusion modules. It also introduces FHAPD, the first large-scale VHR agricultural parcel dataset in China, to support comprehensive evaluation. HBGNet outperforms eight existing methods across multiple datasets (FHAPD, Al4Boundaries, Sentinel-2) in both attribute and geometric metrics, achieving up to 7.5.
Figure 11: Regional Segmentation Performance Across Models
C.4 Unsupervised Domain Adaptation Methods
C.4.1 FDA
FDA[41] is a simple yet effective UDA method that reduces domain gaps by aligning the low-frequency components of source and target images in the Fourier domain. The core idea is that low-frequency information in images (e.g., color and style) mainly accounts for the domain shift, while high-frequency components capture structural details relevant to semantic content. By replacing the low-frequency spectrum of source images with that of target images, FDA transfers global appearance cues without altering the semantic layout. This process improves feature-level alignment and enhances segmentation performance in the target domain without requiring architectural changes or additional supervision.
C.4.2 DAFormer
DAFormer[15] introduces a Transformer-based architecture for UDA in semantic segmentation, leveraging the superior representation capability of vision Transformers. The model integrates a Transformer encoder with a multi-level, context-aware decoder and employs three essential training strategies: rare class sampling to improve pseudo-labels, ImageNet feature distance regularization to stabilize transfer, and learning rate warmup to reduce overfitting. These components collectively enable DAFormer to effectively adapt to the target domain and learn rare or hard classes. DAFormer achieves significant improvements over previous UDA methods, setting new benchmarks on synthetic-to-real segmentation tasks.
C.4.3 HRDA
HRDA[16] proposes a multi-resolution UDA framework to balance fine-detail preservation and long-range context modeling in semantic segmentation. It combines small high-resolution image crops—effective for boundary and small object delineation—with large low-resolution crops that retain broader contextual information. A learned scale attention module adaptively fuses these dual-resolution representations. HRDA significantly improves segmentation quality on domain shift benchmarks while maintaining computational efficiency, and is particularly effective for detailed structures under limited GPU memory.
C.4.4 PiPa
PiPa[23] introduces a unified pixel- and patch-level self-supervised learning framework for UDA in semantic segmentation. Unlike traditional UDA methods that only minimize inter-domain discrepancies, PiPa focuses on enhancing intra-domain consistency by modeling pixel-level intra-class compactness and patch-level context-invariant semantics. This dual-level supervision enables the model to learn more robust and domain-invariant representations. PiPa achieves competitive results on standard UDA benchmarks and can be flexibly combined with other UDA methods to further improve performance without increasing model complexity.
Figure 13: Qualitative edge detection comparisons between PiDiNet [30] and UEAD [53].
Appendix D Evaluation metrics
D.1 Pixel-level evaluation metrics
$Prec.=\frac{TP}{TP+FP}\quad$ (1)
$Rec.=\frac{TP}{TP+FN}\quad$ (2)
$IoU=\frac{TP}{TP+FP+FN}\quad$ (3)
Figure 14: Qualitative edge detection comparisons between REAUNet-Sober [21] and MuGE [52].
$OA=\frac{TP+TN}{TP+FP+FN+TN}\quad$ (4)
$F1=\frac{2\cdot Prec.\times Rec.}{Prec.+Rec.}\quad$ (5)
[…]
D.2 Object-level geometric metrics
$\text{OC}(S_{i})=1-\frac{\text{area}(S_{i}\cap O_{i})}{\text{area}(O_{i})},$ (6)
$\text{UC}(S_{i})=1-\frac{\text{area}(S_{i}\cap O_{i})}{\text{area}(S_{i})},$ (8)
$\text{TC}(S_{i})=\sqrt{\frac{\text{OC}(S_{i})^{2}+\text{UC}(S_{i})^{2}}{2}},$ (10)
D.3 Edge detection evaluation metrics
$F_{t}=\frac{2\cdot P_{t}\cdot R_{t}}{P_{t}+R_{t}}.$ (12)
$\text{AP}=\int_{0}^{1}P(R),dR,$ (15)
Appendix E More results
E.1 More results on Semantic Segmentation
E.1.1 Discussions the results for different regions
E.1.2 More visualization cases
Fig. 12 shows error maps for each model on the same test region, where black denotes true negatives, white true positives, blue false negatives, and red false positives. This visualization highlights that transformer-based methods such as Mask2Former [8] produce fewer false negatives in complex boundary areas. However, overall, the models struggle to accurately segment very fine-grained background details.
E.2 More visualization results on edge detection
Fig. 13 and Fig. 14 present qualitative comparisons of edge detection results on representative regions using four models: UEAD [53], PiDiNet [30], MuGE [52], and REAUNet-Sober [21]. From the visualizations, we observe that UEAD and MuGE tend to produce wider and blurred boundaries, indicating over-smoothed predictions that reduce localization precision. This issue is especially pronounced in densely terraced landscapes, where precise boundary localization is critical. In contrast, PiDiNet and REAUNet-Sober generate sharper and more compact edges, with REAUNet-Sober showing the best alignment with ground truth boundaries across complex spatial structures.
E.4 Visualization cases on domain adaptation
Fig. 16, Fig. 17, and Fig. 18 illustrate the segmentation results of four UDA methods (FDA, DAFormer, HRDA, and PiPa) on six representative domain transfer tasks, with transfers targeting the North (N), South (S), and Global (G) domains, respectively. The results show that HRDA and PiPa generally produce finer boundaries and better preserve parcel structures under domain shift.
Appendix F Limitations and future work
Although GTPBD encompasses major terraced regions worldwide, the current collection is limited to six well-known terraced areas in countries where terracing is concentrated. Some atypical terraced regions in other countries are not included due to the sparse distribution of terraces in those locations. The dataset currently consists only of optical images and does not include other remote sensing modalities, such as multispectral or infrared images.
In future work, the dataset will be expanded by incorporating additional atypical terrace samples and broadening global coverage. Furthermore, GTPBD will be supplemented with multimodal geographical data, including information from Digital Elevation Models (DEMs), slope gradients, spatiotemporal features, and even will be incorporated into Q&\&&A pairs to serve multimodal large language model. Given that the dataset already covers the primary terrace types, future task may focus on domain generalization to identify new terrace types.

### v3

@@ 4.4 Edge Detection Results
Conventional edge detection datasets typically include multi-scale, manually annotated boundaries. To simulate this for the GTPBD dataset, we apply morphological operations (erosion and dilation) to generate edge labels with widths of 1 to 5 pixels, mimicking human-labeled multi-resolution boundaries. Table 3 presents the results of four edge detection models. Among them, REAUNet-Sober achieves the best performance across all three metrics—ODS (65.06%), OIS (63.73%), and AP (70.09%)—significantly outperforming both general-purpose (e.g., PiDiNet), demonstrating its effectiveness in delineating parcel edges under noisy conditions. Visual examples in Fig. 7 (d) further illustrate model robustness, where false positives are marked in red and missed detections in blue.More visualization results and the performance of regions could be found in AppendixE.2
@@
U-Net unet 74.34 54.29 45.72 75.39 62.75 43.57 37.04 40.43
PSPNet pspnet 68.55 71.50 53.84 76.59 69.99 22.66 49.90 39.75
SegFormer segformer 74.71 68.23 55.43 79.05 71.32 27.15 41.17 35.84
Mask2Former mask2former 71.48 73.42 56.79 78.66 72.44 22.04 45.15 35.53
SEANet SEANet 70.69 79.56 59.89 76.07 74.04 26.80 54.70 44.08
REAUNet ReauNet-Sober 73.03 78.01 60.56 80.60 75.44 27.02 42.25 36.07
HBGNet FHAPD/HBGNet 72.50 81.81 62.44 81.20 76.88 27.40 42.52 35.79
@@
S $\rightarrow$ N Source Only 61.57 68.75 48.11 71.07 64.96 61.55 68.58 48.01 71.03 64.88 25.49 49.12 39.13
FDAFDA 75.60 46.72 40.60 73.33 57.75 75.59 46.41 40.36 73.25 57.51 53.95 34.20 45.16
PiPa PiPa 62.74 84.71 56.35 74.41 72.09 62.72 84.05 56.05 74.29 71.84 13.79 54.71 40.91
HRDA HRDA 81.48 59.31 52.26 78.87 68.65 81.56 58.78 51.88 78.73 68.32 42.02 26.27 35.25
DAFormer DAFormer 82.12 58.18 51.64 78.74 68.11 82.15 57.79 51.35 78.64 67.85 41.08 29.49 35.78
S $\rightarrow$ G Source Only 66.26 72.53 52.97 77.63 69.25 66.27 72.46 52.94 77.63 69.23 26.62 48.39 39.95
FDAFDA 67.03 53.79 42.54 74.76 59.68 67.21 52.71 41.93 74.64 59.08 48.20 44.08 46.80
PiPa PiPa 61.33 80.99 53.61 75.66 69.80 61.54 79.33 53.03 75.60 69.31 18.06 54.68 40.71
HRDA HRDA 62.85 73.72 51.35 75.74 67.85 63.15 72.01 50.70 75.69 67.29 22.48 51.45 39.70
DAFormer DAFormer 66.24 62.43 47.36 75.90 64.28 66.43 61.36 46.84 75.81 63.79 38.18 42.73 40.51
N $\rightarrow$ S Source Only 82.18 68.18 59.40 81.21 74.53 82.36 68.03 59.38 81.23 74.51 32.40 33.53 33.46
FDAFDA 71.37 62.66 50.07 74.81 66.73 71.62 62.05 49.80 74.78 66.49 40.43 50.03 45.48
PiPa PiPa 79.28 80.71 66.65 83.72 79.99 79.55 79.91 66.29 83.62 79.73 23.88 39.95 32.91
HRDA HRDA 77.19 84.01 67.30 83.74 80.46 77.51 83.13 66.98 83.77 80.22 19.23 41.83 32.55
DAFormer DAFormer 78.26 82.41 67.05 83.67 80.28 78.59 81.56 66.73 83.60 80.05 19.39 42.21 32.85
N $\rightarrow$ G Source Only 73.83 56.90 47.36 78.03 64.27 73.86 56.86 47.33 78.03 64.25 45.92 33.78 40.30
FDAFDA 66.10 51.38 40.66 73.96 57.82 66.22 50.68 40.27 73.89 57.42 56.64 48.15 52.57
PiPa PiPa 69.60 80.67 59.65 81.05 74.72 69.85 79.30 59.08 80.93 74.28 24.34 42.21 34.45
HRDA HRDA 70.66 83.45 61.98 82.22 76.52 70.97 81.90 61.35 82.08 76.04 20.67 43.05 32.37
DAFormer DAFormer 74.96 67.77 55.26 80.95 71.19 75.21 66.54 54.57 80.76 70.61 32.33 33.14 32.74
G $\rightarrow$ S Source Only 80.06 72.26 61.24 81.56 75.96 80.22 72.10 61.22 81.58 75.94 28.63 37.75 33.50
FDAFDA 66.15 51.92 41.02 69.90 58.18 66.37 51.42 40.79 69.91 57.95 46.02 49.66 47.88
PiPa PiPa 78.06 77.84 63.86 82.23 77.95 77.96 76.65 62.99 81.84 77.30 22.67 41.10 33.29
HRDA HRDA 74.59 76.28 60.54 79.96 75.42 74.90 75.45 60.23 79.91 75.18 25.91 42.75 35.35
DAFormer DAFormer 78.72 76.75 63.56 82.26 77.72 79.01 76.01 63.23 82.18 77.48 26.46 40.55 34.04
G $\rightarrow$ N Source Only 77.03 69.08 57.28 79.90 72.84 77.03 69.03 57.25 79.89 72.81 32.14 35.97 34.10
FDAFDA 53.77 75.31 45.72 65.12 62.75 53.72 74.87 45.51 65.04 62.56 25.75 56.99 44.22
PiPa PiPa 62.13 88.77 57.60 74.51 73.10 62.12 88.10 57.31 74.40 72.86 10.14 54.53 39.22
HRDA HRDA 58.89 83.01 52.56 70.77 68.90 58.85 82.40 52.28 70.66 68.66 20.83 51.16 39.01
DAFormer DAFormer 70.80 72.87 56.03 77.69 71.82 70.81 72.26 55.68 77.56 71.53 28.16 41.19 35.28
Avg Source Only 73.49 68.12 54.39 77.23 70.51 73.53 68.01 54.29 77.24 70.44 31.70 42.59 36.74
FDAFDA 70.00 56.13 43.25 71.99 62.22 70.22 55.94 42.92 71.92 61.91 45.17 47.19 47.10
PiPa PiPa 69.02 82.12 59.62 78.60 74.38 69.29 81.22 59.13 78.46 74.07 17.15 47.88 36.91
HRDA HRDA 70.78 76.63 58.50 78.45 73.30 71.16 75.61 57.91 78.34 72.95 25.19 42.75 35.93
DAFormer DAFormer 75.48 69.90 56.82 79.72 72.40 75.63 69.25 56.23 79.58 72.22 31.10 38.22 36.20
@@ 4.5 Boundary-integration terraced parcel extraction results
Building on the boundary constraint theory proposed by Yuan et al. Yuhui2020, which highlights the importance of edge orientation and intensity in enhancing object-level delineation, we explore boundary-integration strategies for agricultural parcel segmentation. We compare four general-purpose segmentation models (U-Net, PSPNet, SegFormer, and Mask2Former) with three agriculture-specific ones (HBGNet, REAUNet, and SEANet). To incorporate edge cues, we introduce a boundary detection pipeline (see Fig. 7) that explicitly integrates edge information into the segmentation process. REAUNet-Sober achieves the best performance among the trained edge detectors, and its binarized output is used to guide segmentation refinement. As shown in Table 4, Mask2Former delivers the best object-level performance, achieving the lowest GOC (22.04%) and GTC (35.53%), despite lower pixel-level scores. In contrast, HBGNet attains the highest pixel-level accuracy (Recall: 81.81%, IoU: 62.4%, OA: 82.20%, F1: 76.88%) but suffers from higher object-level errors, indicating a trade-off between pixel precision and structural integrity.
@@ 4.6 Unsupervised domain adaptation results
We evaluate domain robustness of segmentation models through UDA experiments across three distinct domains in GTPBD (see Sec. 3.3). The Source only and other four domain adaptation methods are benchmarked: FDA, DAFormer, HRDA, and PiPa. As shown in Table 5, across all six domain adaptation directions, PiPa achieves the best performance in pixel-level metrics (Rec., IoU, F1), demonstrating stable cross-domain generalization. Notably, HRDA performs best when the source domain is N, achieving superior results across nearly all metrics under this setting. After adding boundary results (integrated by the performance of REAUNet-Sober), further achieves the lowest GOC and GTC, indicating a stronger ability to preserve parcel boundaries and maintain object completeness, while DAFormer attains the lowest GUC, reflecting its advantage in reducing under-segmentation and improving boundary coherence.Visualization comparisons are provided in Appendix E.4.
@@ 5 Conclusion
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant T2125006 and Grant 42401415; in part by the Shenzhen Science and Technology Program under Grant KCXFZ20240903093759004 and Grant KJZD20230923115106012; in part by the Fundamental Research Funds for the Central Universities, Sun Yat-sen University, under Project 24xkjc002; and in part by the Jiangsu Innovation Capacity Building Program under Project BM2022028.